### v1

# Generalisation in Multi-task Offline Q-Learning

Abstract
We study offline multitask reinforcement learning in settings where multiple tasks share a low-rank representation of their action-value functions. In this regime, a learner is provided with fixed datasets collected from several related tasks, without access to further online interaction, and seeks to exploit shared structure to improve statistical efficiency and generalization. We analyze a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific value functions via Bellman error minimization on offline data. Under standard realizability and coverage assumptions commonly used in offline reinforcement learning, we establish finite-sample generalization guarantees for the learned value functions. Our analysis explicitly characterizes how pooling data across tasks improves estimation accuracy, yielding a $1/\sqrt{nT}$ dependence on the total number of samples across tasks, while retaining the usual dependence on the horizon and concentrability coefficients arising from distribution shift. In addition, we consider a downstream offline setting in which a new task shares the same underlying representation as the upstream tasks. We study how reusing the representation learned during the multitask phase affects value estimation for this new task, and show that it can reduce the effective complexity of downstream learning relative to learning from scratch. Together, our results clarify the role of shared representations in multitask offline Q-learning and provide theoretical insight into when and how multitask structure can improve generalization in model-free, value-based reinforcement learning.
Keywords Reinforcement Learning $\cdot$ Statistical Machine Learning $\cdot$ Representation Learning $\cdot$ Multi-task Learning
1 Introduction
In many real-world reinforcement learning (RL) applications, collecting new experience through online interaction is costly, time-consuming, or unsafe. Offline reinforcement learning addresses this challenge by learning decision-making policies from fixed datasets collected by previously deployed behavior policies, without requiring further interaction with the environment. In practice, such datasets often arise from multiple related tasks, motivating the study of offline multitask reinforcement learning, where a learner seeks to exploit shared structure across tasks to improve generalization and sample efficiency [10, 15, 11].
A common assumption in multitask RL is that tasks share an underlying low-dimensional representation, even though their rewards or dynamics may differ. Learning such shared representations can substantially reduce statistical complexity, particularly when the number of samples per task is limited. Empirically, representation sharing has been shown to improve transfer and generalization across tasks, but providing theoretical guarantees in the offline setting remains challenging due to distribution shift, limited coverage, and recursive error propagation through Bellman updates.
Most existing theoretical analyses of multitask representation learning in RL focus on model-based approaches, where shared structure is recovered through maximum-likelihood estimation of transition models or latent dynamics. In contrast, many practical algorithms are model-free and value-based, relying on fitted Q-iteration and Bellman error minimization. Despite their widespread use, theoretical guarantees for such methods in multitask, offline settings are comparatively limited.
In this work, we study offline multitask reinforcement learning through the lens of value-based methods. We analyze a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific value functions from offline data via Bellman error minimization. We adopt a low-rank realizability assumption on the optimal action-value functions, where tasks share a common (unknown) representation but have task-dependent linear decoders [11]. Within this framework, we provide finite-sample generalization guarantees that explicitly quantify how pooling data across $T$ tasks improves statistical efficiency, yielding error rates that scale as $O(1/\sqrt{nT})$ in the number of samples per task $n$ and the number of tasks $T$, up to standard horizon- and coverage-dependent factors.
Beyond the upstream multitask setting, we also consider a downstream offline scenario in which a new task shares the same underlying representation as the source tasks. We analyze how reusing the representation learned during multitask training affects value estimation for the new task, and show that it can reduce the effective complexity of downstream learning compared to learning the representation from scratch.
Contributions. Our main contributions are as follows:
  • We provide a finite-sample analysis of multitask fitted Q-iteration in the offline, model-free setting under a shared low-rank representation assumption.
  • Our bounds explicitly capture the statistical benefits of multitask data pooling, with estimation error scaling as $O(1/\sqrt{nT})$ under standard realizability and coverage assumptions.
  • We characterize how Bellman error propagates across the horizon in the multitask setting, highlighting the role of concentrability in offline value-based learning.
  • We analyze a downstream offline task and show that reusing representations learned from upstream tasks can improve sample efficiency relative to independent learning.
2 Related Work
Offline Reinforcement Learning.
Offline or batch reinforcement learning studies policy learning from fixed datasets without further interaction with the environment. Classical approaches rely on fitted Q-iteration [8, 21] and subsequent deep extensions such as BCQ, CQL, and IQL [9, 14, 13]. Theoretical analyses of offline RL have focused on characterizing sample complexity under concentrability assumptions [19, 2, 12]. Our work extends this line of research to the multitask setting, where shared structure across tasks is exploited to improve statistical efficiency in the offline regime.
Multitask and Representation Learning in RL.
Representation learning for reinforcement learning has been studied as a means to transfer information across related tasks [3, 23, 16]. Recent theoretical progress has formalized multitask representation learning as a low-rank factorization of value or transition functions [1, 7]. These works typically assume access to interactive data or known models. Our formulation builds directly on this multitask structure but focuses on offline data settings and on value-based (Q-function) estimation rather than model estimation.
Model-Based Multitask RL.
Most existing theoretical treatments of multitask offline RL are model-based. The closest to our work is Domingues et al. [6], who propose an offline multitask framework that estimates shared low-rank transition models via maximum likelihood estimation (MLE). Their analysis provides generalization guarantees for multitask model learning but does not extend to model-free Q-learning. Our approach replaces the MLE oracle with a Bellman-error minimization oracle, leading to a model-free formulation and analysis that directly connects to practical Q-learning methods.
Model-Free and Value-Based Theoretical Analyses.
Model-free RL methods, such as Q-learning and actor–critic algorithms, are known to suffer from instability when trained from offline data [17]. Recent theoretical results for single-task offline Q-learning [12, 20] establish finite-sample guarantees via Bellman error concentration. Our work extends these ideas to the multitask setting and develops new generalization bounds for shared representation learning under mean-squared Bellman error minimization, thereby unifying multitask representation learning and offline value-based RL.
Summary.
In summary, prior work on multitask reinforcement learning has largely focused on model-based estimation and online interaction, while existing analyses of offline Q-learning are predominantly single-task. Our work complements these directions by studying model-free, value-based learning in the multitask offline setting and providing finite-sample generalization guarantees that parallel those obtained in model-based multitask analyses under comparable assumptions.
3 Preliminary
Episodic MDP. An episodic discrete-time Markov Decision Process (MDP), represented by $\mathcal{M}=(\mathcal{S},\mathcal{A},H,P,r)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is a finite action space. $K$ is defined as the cardinality of $\mathcal{A}$, and $H$ is the finite episode length, $P=\{P_{h}\}_{h\in H}$ are the state transition probability distributions with $P_{h}:\mathcal{S\times A}\to\Delta(\mathcal{S})$, and $r=\{r_{h}\}_{h\in[H]}$ are the deterministic reward functions with $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$. Following (Jiang et al., 2017; Sun et al., 2019), we assume that the initial state $s_{1}$ is fixed for each episode. A policy $\pi$ is a collection of $H$ functions $\{\pi_{h}:\mathcal{S}\to\mathcal{A}\}_{h\in[H]}$ where $\pi_{h}(s)$ is the action that the agent takes at state $s$ and at the $h$-th step in the episode. Given a starting state $s_{h}$, $s_{h^{\prime}}\sim(P,\pi)$ denotes a state sampled by executing policy $\pi$ under the transition model $P$ for $h^{\prime}-h$ steps and $\mathbb{E}_{(s_{h},a_{h})\sim(P,\pi)}[\cdot]$ denotes the expectation over states $s_{h}\sim(P,\pi)$ and actions $a_{h}\sim\pi$. Moreover, for each $h\in[H]$, we define the value function under policy $\pi$ when starting from an arbitrary state $s_{h}=s$ at the $h$-th time step as
$V^{\pi}_{h,P,r}(s)=\mathbb{E}_{(s_{h}^{\prime},a_{h}^{\prime})\sim(P,\pi)}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})|s_{h}=s\right]$
We define the action-value function, also called the $Q$ function, for a given state action pair, $(s,a)$ as
$Q^{\pi}_{h,P,r}(s,a)=\mathbb{E}_{(s_{h}^{\prime},a_{h}^{\prime})\sim(P,\pi)}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})|s_{h}=s,a_{h}=a\right]$
Multitask Episodic MDPs.
This form of learning arises from the fact that may agents would need to perform multiple tasks in the same environment, which is determined by a common state and action spaces $\mathcal{S\times A}$. Each task $t\in[T]$ (from a set of $T$ tasks) can be represented by an episodic discrete time MDP: $(\mathcal{S},\mathcal{A},H,P^{(t)},r^{(t)})$. We note that the MDPs for each task share the same state space, action space and episode lengths, while differing in the transition probabilities and rewards. A policy $\pi$ in a multitask episodic MDP can also be redefined as a set of policies $\pi=\{\pi^{(t)}\}_{t\in[T]}$, such that $\pi^{(t)}$ is itself a collection $\{\pi_{h}^{(t)}:\mathcal{S}\to\mathcal{A}\}_{h\in[H]}$. Based on this we can also now use task specific value function as follows:
$V^{\pi,t}_{h,P,r}(s)=\mathbb{E}_{(s_{h}^{\prime},a_{h}^{\prime})\sim(P,\pi^{(t)})}\left[\sum_{h^{\prime}=h}^{H}r^{(t)}_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})|s_{h}=s\right]$
And similarly we also have a task specific Q function:
$Q^{\pi,t}_{h,P,r}(s,a)=\mathbb{E}_{(s_{h}^{\prime},a_{h}^{\prime})\sim(P,\pi^{(t)})}\left[\sum_{h^{\prime}=h}^{H}r^{(t)}_{h^{\prime}}(s_{h^{\prime}},a_{h^{\prime}})|s_{h}=s,a_{h}=a\right]$
3.1 Multitask Offline Q-Learning with Shared Representations
In multitask offline Q-learning, the agent is provided with offline datasets collected from $T$ related source tasks, denoted $\{\mathcal{D}^{(t)}\}_{t\in[T]}$. Each task $t$ is associated with an episodic Markov decision process $\mathcal{M}^{(t)}=(\mathcal{S},\mathcal{A},H,P^{(*,t)},r^{(*,t)})$. All tasks share a common state–action space and horizon $H$, but may differ in their transition dynamics $P^{(*,t)}$ or reward functions $r^{(*,t)}$. We assume that the tasks share a latent representation of the state space: there exists a feature encoder $\phi^{*}:\mathcal{S\times A}\to\mathbb{R}^{d}$ such that the optimal action-value functions satisfy
$Q_{h}^{(*,t)}(s,a)=\langle\phi^{*}(s,a),\omega_{h}^{(*,t)}\rangle$
for some task-specific decoders: $\omega_{h}^{(*,t)}$. This shared structure allows the agent to transfer information across tasks through a common embedding $\phi^{*}$. We adopt this low-rank structure as it is popular and has been emperically justified in previous work ([4]). The primary theoretical benefit of using a linear decoder is its minimal contribution to the overall function class capacity, which is crucial because this capacity scales rapidly demonstrated in Theorem 1.
The learning algorithm has access only to offline datasets
$\mathcal{D}^{(t)}=\bigcup_{h\in[H]}\mathcal{D}^{(t)}_{h},\quad\mathcal{D}^{(t)}_{h}=\{(s_{h}^{(i,t)},a_{h}^{(i,t)},r_{h}^{(i,t)},s_{h+1}^{(i,t)})\}_{i=1}^{n}$
collected under fixed behavior policies $\pi_{b}^{(t)}$ for each task. Each dataset therefore consists of transitions (state and action) $s_{h+1}^{(i,t)},a_{h}^{(i,t)}\sim P^{(*,t)}_{h},\pi_{b}^{(t)}$ and observed rewards $r_{h}^{(i,t)}$. The goal of upstream multitask learning is to learn a shared representation $\hat{\phi}$ and task-specific heads $\{\hat{\psi}^{(t)}\}_{t=1}^{T}$ that minimize the expected Q-Bellman error across tasks using only the offline data. Formally, at each step $h$, we minimise the following Q-bellman error:
$\hat{\mathcal{L}}(\phi,\{\psi^{(t)}\}_{t\in[T]})=\frac{1}{Tn}\sum_{t=1}^{T}\sum_{i=1}^{n}\Big(\langle\phi(s_{h}^{(i,t)},a_{h}^{(i,t)}),w_{h}^{(t)}\rangle-r_{h}^{(i,t)}-\max_{a^{\prime}}\langle\phi(s_{h+1}^{(i,t)},a^{\prime}),w_{h+1}^{(t)}\rangle)\Big)^{2}$
3.2 Role of the Concentrability Coefficient
A central difficulty in offline reinforcement learning is distribution shift between the data-generating distribution and the distributions induced by Bellman backups. Since learning is performed using samples from a fixed behavior policy, errors measured under the behavior distribution may be amplified when propagated backward through the Bellman operator.
Let $\mu_{b}$ denote the state–action distribution induced by the behavior policy. For any policy $\pi$ and stage $h$, let $\mu_{h}^{\pi}$ denote the state–action distribution induced at stage $h$ by following $\pi$. We quantify distribution shift using a concentrability coefficient
$\lambda_{\max}\;\triangleq\;\sup_{h\in[H]}\sup_{\pi}\left\|\frac{\mu_{h}^{\pi}}{\mu_{b}}\right\|_{\infty},$
which upper bounds the density ratio between Bellman-induced distributions and the offline data distribution.
In the online or on-policy setting, Bellman operators are contractive in suitable norms, leading to controlled error propagation. In contrast, in the offline setting, contraction may fail under $\mu_{b}$, and $\lambda_{\max}$ appears as a multiplicative factor in generalization bounds. Consequently, the bounds in Theorems 1 and 2 scale polynomially with $\lambda_{\max}$, reflecting the inherent difficulty of offline value-based learning under distribution shift. When the behavior policy provides sufficient coverage, $\lambda_{\max}$ remains bounded, yielding meaningful finite-sample guarantees.
This multitask empirical-risk objective performs joint representation learning and value-function regression across all tasks, providing a shared encoder that generalizes to unseen downstream environments. To ensure bounded sample complexity, we adopt the coverage condition standard in offline RL, assuming that each behavior policy $\pi_{b}^{(t)}$ sufficiently covers the state–action distribution induced by optimal or target policies. Under this assumption, our analysis establishes finite-sample generalization guarantees for the learned representation $\hat{\phi}$ and task-specific Q-functions $\hat{Q}^{(t)}$, showing that shared representations learned via multitask Q-learning achieve comparable sample efficiency to multitask model-based approaches. In Section 4.2, we formalize this result and provide our main finite-sample bound, demonstrating that the proposed multitask Q-learning framework admits strong theoretical guarantees in the offline setting.
Algorithm 1 Multi-task Fitted Q-Iteration
[…]
2:# datasets for each task $t$
3:Initialize parameters $\Theta\leftarrow\Theta_{0}$, iteration counter $k\leftarrow 0$
4:for $h=H$ down to 11 do
5:  repeat
6:   Compute targets:
7:   for $t=1$ to $T$ do
8:     $\mathcal{Y}_{t}^{(k+1)}\leftarrow\left\{R_{t}(s,a)+\gamma\max_{a^{\prime}}Q^{(k)}(s^{\prime},a^{\prime})\;\middle|\;(s,a,s^{\prime})\in\mathcal{D}_{t}\right\}$
9:   end for
10:   Multi-task update:
11:   $\Theta^{(k+1)}\leftarrow\operatorname{MTL}\!\left(\bigcup_{t=1}^{T}\mathcal{D}_{t},\;\bigcup_{t=1}^{T}\mathcal{Y}_{t}^{(k)}\right)$
12:   $d\Theta\leftarrow\lVert\Theta^{(k+1)}-\Theta^{(k)}\rVert$
13:   $k\leftarrow k+1$
14:  until $d\Theta<\epsilon$ or $k\geq\text{MaxIter}$
15:end for
16:return $\Theta=\{\theta_{t}\}_{t=1}^{T}$, where $Q_{t}(s,a)=f_{\theta_{t}}(s,a)$
4 Upstream Offline Multitask Representation Learning
In this section, we introduce our offline Multitask Q-Learning designed for upstream offline multitask RL in a low-rank Q-learner setting and describe its theoretical properties.
4.1 Multitask Fitted Q-Iteration Algorithm
The details of the algorithm MQRL (Multitask Q-Learning for Representation Learning) are depicted in Algorithm 1. The agent receives all offline datasets $\{\mathcal{D}^{(t)}\}_{t\in[T]}$ and jointly estimates the shared encoder $\phi_{h}$ and the task-specific decoders $\{\psi_{h}^{(t)}\}_{t\in[T]}$ by minimizing the empirical Bellman error across all tasks at each stage $h$:
$(\hat{\phi}_{h},\hat{\psi}_{h}^{(1)},\dots,\hat{\psi}_{h}^{(T)})=\arg\min_{\phi_{h}\in\Phi,\psi_{h}^{(1)},\dots,\psi_{h}^{(T)}\in\Psi}\hat{\mathcal{L}}(\phi_{h},\{\psi_{t}\}_{t\in[T]})$ (1)
This objective acts as a multitask empirical-risk minimization oracle, serving as the offline Q-learning analogue of the maximum-likelihood oracle used in prior multitask model-based work (Agarwal et al., 2020b; Domingues et al., 2021). In practice, the optimization in 1 can be efficiently approximated whenever the function classes $\Phi$ and $\Psi$ are parameterized by neural networks or other differentiable architectures, allowing gradient-based minimization. The shared encoder $\phi_{h}$ aggregates statistical information across all tasks, thereby promoting representation sharing, while each decoder $\psi_{h}^{(t)}$ specializes to its respective task to approximate its local value function. In Algorithm 1 we prefer $\Theta$ to represent the combination of $\phi$ and $\omega_{h}^{(t)}$.
4.2 Theoretical Results
To facilitate the analysis, we adopt a realizability assumption that is standard in the low-rank Q-learning literature.prog
Assumption 3.1 (Low-rank realizability). There exists a shared representation $\phi^{*}\in\Phi$ and task- and stage-specific weight vectors $w_{h}^{(*,t)}\in\Psi$ such that, for all $h\in[H]$ and $t\in[T]$,
$Q_{h}^{*,t}(s,a)=\langle\phi^{*}(s,a),w_{h}^{(*,t)}\rangle.$
We assume $\|\phi(s,a)\|_{2}\leq 1$ for all $\phi\in\Phi$ and $\|w\|_{2}\leq 1$ for all $w\in\Psi$.
For simplicity assume that $\Phi$ and $\Psi$ are finite sets. The main theorem here is the following:
Theorem 1.
Under the low-rank MDP structure (Assumption 3.1), we establish finite-sample generalization bounds for multitask offline Q-learning.
  • (a)
    Under realizability, with probability at least $1-\delta$, for any $h\in[H]$,
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{(s_{h},a_{h})\sim\mu_{b}^{(t)}}\left[\|\hat{Q}_{h}^{t}-Q^{\pi_{b}^{t}}_{h}\|^{2}\right]\leq B\sqrt{\frac{2\log(2|\Phi||\Psi|^{T}H/\delta)}{nT}}.$
    This bound exhibits the expected $O(1/\sqrt{nT})$ statistical scaling due to pooling data across tasks.
  • (b)
    With probability $1-\delta$, for any $h\in[H-1]$,
[…]
    $\quad\quad+\ \sqrt{\mathcal{O}\left(\frac{2B}{3}\log\left(\frac{2|\Phi||\Psi|^{T}}{\delta}\right)+\sqrt{\frac{4B^{2}}{9}\log^{2}\left(\frac{2|\Phi||\Psi|^{T}}{\delta}\right)+8\sigma^{2}\log\left(\frac{2|\Phi||\Psi|^{T}}{\delta}\right)}\right)}$
  • (c)
    Averaging across tasks, with high probability,
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{\mu_{b}^{(t)}}\left[\|\hat{Q}_{1}^{t}-Q_{1}^{*,t}\|_{L_{2}(\mu_{b})}\right]\lesssim\tilde{O}\left(H\mathcal{E}_{\text{Approx}}+H^{2}\lambda_{\max,\sup}\sqrt{\frac{|\Phi||\Psi|^{T}}{nT}}+H^{3}\lambda_{\max,\sup}\frac{|\Phi||\Psi|^{T}}{nT}\right),$
    where $\mathcal{E}_{\text{Approx}}$ bounds the layer-wise approximation error, and $\lambda_{\max,\sup}$ is the supremum concentrability coefficient handling distribution shift. This final bound unrolls the recursion in (b), yielding polynomial dependence on the horizon $H$.
Remarks.
Part (a) provides a uniform generalization bound for the fitted Q-function at a fixed stage, highlighting the $1/\sqrt{nT}$ improvement obtained by pooling data across tasks. Part (b) characterizes how estimation error propagates backward through Bellman updates in the offline setting, with the concentrability coefficient $\lambda_{\max}$ accounting for distribution shift. This also Part (c) unrolls this recursion across the horizon to obtain a bound on the initial-stage error, with polynomial dependence on $H$ and explicit dependence on the multitask sample size. Complete proofs are deferred to the appendix.
5 Generalization Bounds for Infinite Hypothesis Sets
When the hypothesis spaces for the encoder $\Phi$ and the decoders $\Psi$ are infinite, such as those parameterized by continuous function approximators, the Union Bound approach utilized in Theorem 1 is no longer applicable. We must instead rely on uniform convergence theory based on the Rademacher complexity to quantify the capacity of the Multitask Hypothesis Space $\mathcal{F}=\Phi\times\Psi^{T}$.
5.1 Multitask Loss Class and Rademacher Complexity
The generalization gap is determined by the complexity of the class of functions that measure the squared Bellman residual. We define the multitask loss function class, $\mathcal{G}$, as:
$\mathcal{G}=\left\{g_{f}:f\in\mathcal{F}\right\}$
where $g_{f}$ is the squared Bellman residual for a function $f=(\phi,\{\psi^{(t)}\})\in\Phi\times\Psi^{T}$ at step $h$:
$g_{f}(s,a,y)=\frac{1}{T}\sum_{t=1}^{T}\left(Q_{h}^{t}(s,a;f)-\text{y}\right)^{2}$
Here, $y=T_{h}^{*}\hat{Q}_{h+1}^{t}$ is the empirical Bellman target used for the regression. I would like the reader to refer to the appendix section for a detailed explanation of Raedmacher complexity and the associated analysis and proof.
Definition 1 (Expected Rademacher Complexity).
The Expected Rademacher Complexity of the loss class $\mathcal{G}$, denoted $\mathcal{R}(\mathcal{G})$, measures the ability of $\mathcal{G}$ to fit random noise on the training data:
$\mathcal{R}(\mathcal{G})=\mathbb{E}_{\mathcal{D},\mathbf{\sigma}}\left[\sup_{g\in\mathcal{G}}\frac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\sigma_{i}^{(t)}g(s_{h}^{(i,t)},a_{h}^{(i,t)})\right]$
where $\mathbf{\sigma}=\{\sigma_{i}^{(t)}\}_{i=1,t=1}^{n,T}$ are independent Rademacher random variables ($\text{Pr}(\sigma=\pm 1)=1/2$).
5.2 Generalization Bound via Rademacher Complexity
By applying the standard generalization theorem (e.g., using McDiarmid’s inequality and the Rademacher complexity bound), we can replace the cardinality-dependent term in Theorem 3.1 with the complexity measure $\mathcal{R}(\mathcal{G})$.
We define the (average) suboptimality gap at stage $h$ as
$\Delta_{h}\;\triangleq\;\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{(s_{h},a_{h})\sim\mu_{b}^{(t)}}\!\left[\,\|\hat{Q}_{h}^{t}(s_{h},a_{h})-Q_{h}^{*,t}(s_{h},a_{h})\|_{L_{2}(\mu_{b})}\,\right].$
For the downstream task, $\Delta_{h}^{\text{downstream}}$ denotes the corresponding quantity for the single novel task.
Theorem 2.
Consider a downstream offline task that shares the same low-rank representation as the upstream multitask setting. Let $\hat{\phi}$ denote the representation learned from the upstream tasks, and let $\mathcal{G}$ be the corresponding class of task-specific decoders. Then, with probability at least $1-\delta$, the downstream suboptimality gap at the initial stage satisfies
$\Delta_{1}^{\text{downstream}}\;\lesssim\;H\lambda_{\max,\sup}\epsilon_{\text{irred}}^{\text{eff}}\;+\;H^{2}\lambda_{\max,\sup}\mathcal{R}(\mathcal{G})\;+\;H^{3}\lambda_{\max,\sup}\frac{\log(1/\delta)}{n},$
where $\epsilon_{\text{irred}}^{\text{eff}}$ denotes the effective irreducible error induced by using the upstream representation $\hat{\phi}$, $\mathcal{R}(\mathcal{G})$ is the Rademacher complexity of the decoder class $\mathcal{G}$, and $\lambda_{\max,\sup}$ is the supremum concentrability coefficient.
This result shows that, when the representation is fixed using upstream multitask data, downstream learning reduces to estimating a lower-complexity decoder class. In particular, the statistical term depends on the Rademacher complexity $\mathcal{R}(\mathcal{G})$, rather than the complexity of the full hypothesis class used during upstream learning. For linear decoders over bounded representations, $\mathcal{R}(\mathcal{G})=O(1/\sqrt{n})$, yielding improved statistical rates compared to learning the representation from scratch.
6 Experiments
The goal of our experiments is not to demonstrate empirical performance, but to provide controlled sanity checks that illustrate the scaling behavior predicted by the theoretical results. In particular, we focus on how estimation error varies as a function of the number of tasks $T$, the number of samples per task $n$, and the horizon $H$, while holding other parameters fixed.
Experimental setup.
We consider synthetic multitask Markov decision processes constructed to satisfy the low-rank realizability assumptions used in our analysis. Each task shares a common representation of dimension $d$, with task-specific linear decoders. Offline datasets are generated using fixed behavior policies, and learning is performed using multitask fitted Q-iteration with squared Bellman error minimization. Unless otherwise stated, results are averaged over multiple independent runs to reduce stochastic variability.
Scaling with the number of tasks.
We first study how the estimation error varies as the number of tasks $T$ increases while keeping the number of samples per task $n$ and the horizon $H$ fixed. Consistent with Theorem 1, the observed error decreases as $T$ grows, exhibiting the expected $O(1/\sqrt{nT})$ dependence due to multitask data pooling.
Scaling with the number of samples per task.
Next, we vary the number of samples per task $n$ while fixing $T$ and $H$. As predicted by the theory, the estimation error decreases at a rate proportional to $O(1/\sqrt{n})$, reflecting standard statistical scaling in value-based learning.
Scaling with the horizon.
Finally, we examine the dependence of the estimation error on the horizon $H$. We observe polynomial growth in $H$, consistent with the horizon-dependent terms appearing in our theoretical bounds. This behavior highlights the effect of recursive error propagation in offline value-based learning.
We emphasize that these experiments are conducted in simplified settings where the modeling assumptions hold by construction, and are intended solely to illustrate qualitative scaling trends rather than to serve as comprehensive empirical evaluations.
Figure 1: T-Scaling: Proof of Multi-task Efficiency. Bound: $O(1/\sqrt{nT})$. As the number of tasks increases (with fixed $n=200$, $H=5$), the average MSE to the true $Q^{*}$ decreases consistently, confirming the multitask efficiency predicted by our bounds.
Figure 2: n-Scaling: Proof of Sample Consistency. Bound: $O(1/\sqrt{n})$. With fixed $T=5$ and $H=5$, increasing samples per task $n$ leads to a clear downward trend in error, consistent with our statistical rates despite natural training variance.
Figure 3: H-Scaling: Proof of Recursive Error Propagation. Bound: $O(H^{3}C_{\max})$. For fixed $T=5$ and $n=500$, the total statistical error grows smoothly and polynomially with horizon $H$, validating the sharpness of our Bernstein-based recursive analysis over looser alternatives.
When varying the number of tasks $T$ while fixing the number of samples per task $n$ and the horizon $H$, we observe a clear decrease in estimation error as $T$ increases. Since the total amount of offline data grows linearly with $T$ (total samples =nT=nT), Theorem 1(b) predicts a statistical term scaling as $O(1/\sqrt{nT})$, which reduces to $O(1/\sqrt{T})$ in this regime. After unrolling the recursion in Theorem 1(c), the dominant contribution to the suboptimality gap $\Delta_{1}$ scales as $O(H^{2}\lambda_{\max}/\sqrt{nT})$. The observed downward trend in error as $T$ grows is consistent with this predicted behavior and illustrates the benefit of pooling data across tasks when learning a shared representation.
Importantly, the error decreases with increasing $T$, rather than increasing, which would indicate negative transfer or model misspecification. The observed fluctuations are expected due to stochastic optimization and finite-sample effects, but they do not obscure the overall scaling trend.
Next, we fix the number of tasks $T=5$ and the horizon $H=5$, and vary the number of samples per task $n$. In this setting, the total dataset size grows linearly with $n$, and Theorems 1(b) and 1(c) predict a leading statistical term scaling as $O(1/\sqrt{n})$. Consistent with this prediction, the estimation error decreases as $n$ increases, despite noticeable variance across runs. These fluctuations are expected in multitask training with finite $T$ and do not alter the qualitative $1/\sqrt{n}$ dependence.
Finally, when varying the horizon $H$ while holding $T$ and $n$ fixed, we observe polynomial growth in the estimation error. This behavior is consistent with the horizon-dependent terms in Theorem 1(c), and rules out exponential error propagation. Overall, these experiments illustrate that multitask offline Q-learning exhibits the qualitative scaling behavior predicted by the theory with respect to the number of tasks, samples per task, and planning horizon.
7 Conclusion
We studied multitask offline reinforcement learning in settings where tasks share a common low-rank representation of their action-value functions. Focusing on a model-free, value-based approach, we analyzed a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific decoders from offline data. Under standard realizability and coverage assumptions, we derived finite-sample generalization guarantees that explicitly characterize how pooling data across tasks improves statistical efficiency.
A key takeaway from our analysis is that multitask offline Q-learning achieves final suboptimality bounds that are competitive with existing theoretical results for both offline and multitask reinforcement learning. In particular, despite operating in a strictly more challenging setting—combining offline learning, multitask representation sharing, and model-free Bellman regression—our bounds match the best-known polynomial dependence on the horizon and exhibit the expected $1/\sqrt{nT}$ scaling due to multitask data aggregation. This suggests that multitask structure can be exploited effectively without incurring additional statistical penalties relative to single-task or model-based approaches.
We further analyzed a downstream offline learning setting in which a novel task shares the representation learned from upstream multitask data. Our results show that reusing this representation reduces the effective complexity of downstream learning and leads to improved sample efficiency compared to learning from scratch, while preserving the same order of horizon dependence.
Taken together, our findings indicate that multitask offline representation learning via value-based methods is not only theoretically sound, but also statistically competitive with existing approaches. Rather than improving rates through stronger assumptions or algorithmic pessimism, our analysis demonstrates that carefully exploiting shared structure suffices to recover the same order-optimal guarantees in a model-free framework. We view this work as a step toward a unified understanding of multitask and offline reinforcement learning, and an invitation to further investigate lower bounds, alternative coverage conditions, and extensions to richer function approximation classes.
Table 1: Comparison of final suboptimality / estimation error bounds for offline and multitask reinforcement learning methods. Here $T$ denotes the number of tasks, $n$ the number of samples per task, $H$ the horizon, and $\lambda$ a concentrability coefficient. Logarithmic factors are omitted.
Method Setting Final Error Bound
Our bound (MFQI) Offline, multitask $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{nT}}\right)$
MORL ([11] ) Offline, multitask $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{nT}}\right)$
REFUEL ([5]) Offline / online, multitask $\tilde{O}\!\left(\varepsilon_{\text{rep}}+{H^{2}\lambda}/{\sqrt{n}}\right)$
Pessimistic Q-learning ([12]) Offline, single-task $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{n}}\right)$
Appendix A Proof of Theorem 1a
Theorem (1a). [Theorem 1 part (a)] Under Assumption 3.1, with probability at least $1-\delta$, for any step $h\in[H]$, we have:
[…]
where $\hat{\phi},\hat{P}^{(1)},\ldots,\hat{P}^{(T)}$ are the output of Algorithm 1 and an appropriete constant $B$.
Proof. We start with considering the following random variables: $X_{i}$, where for a fixed $f\in\mathcal{F}$, let $X_{1},X_{2},\ldots,X_{n}$ be i.i.ds, where $X_{i}=l(\hat{y}_{i},y_{i})$, we start with the Hoeffding’s inequality:
$\mathbf{Pr}(|\mu-\hat{\mu}|\geq\epsilon)\leq\exp\left(-\frac{2n\epsilon^{2}}{B^{2}}\right)$
Where $B$ is an upper bound on $X_{i}$, $\hat{\mu}_{f}$ is the empirical mean of $X_{i}$s and $\mu_{f}$ is the true mean corresponding to the true distributions. We can use this combined with Union bounds to get the following bound: with probability $1-\delta$, we have:
$\forall f\in\mathcal{F}:\quad\mu_{f}\leq\hat{\mu}_{f}+B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}$
We have $f^{*}=\arg\min_{f\in\mathcal{F}}\mu_{f}$ and $\hat{f}=\arg\min_{f\in\mathcal{F}}\ \hat{\mu}_{f}$. Using the inequality we have developed:
$\mu_{\hat{f}}\leq\hat{\mu}_{\hat{f}}+B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}\leq\hat{\mu}_{f^{*}}+B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}$
Now reapplying the inequality obtained from the union bound for $f=f^{*}$ we get the following:
$\mu_{\hat{f}}\leq\hat{\mu}_{f^{*}}+B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}\leq\mu_{f^{*}}+2B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}$
Note that this is an upper bound on the empirical mean and not the true mean. However, even this stands correct as a different (necessary) bi-product of the standard union bound. Now since $\mu_{f^{*}}=0$, we can simplify the following inequality as follows:
$\mu_{\hat{f}}\leq 2B\sqrt{\frac{\log(2|\mathcal{F}|/\delta)}{2n}}$
Our interpretation of the loss function, sticking to our current case can be interpreted as follows: at the $h$-th step represents the loss function we have as the Q bellman error term, to obtain the expand the following terms as:
$\mu_{\hat{f}}=\frac{1}{T}\sum_{t\in[T]}\mathbb{E}_{(s_{h},a_{h})\sim(P^{(*,t)},\pi^{b}_{t})}\Big[(Q_{\theta,h}^{t}(s_{h},a_{h})-Q_{h}^{\pi_{b}^{t}}(s_{h},a_{h}))^{2}\Big]$
Note that $B$ represents the bound on the $X_{i}$s, $B=Q_{max}^{2}={\Big(\frac{1-\gamma^{H}}{1-\gamma}\Big)}^{2}$, since our rewards are bounded, which can be treated as a constant for now. Putting this together, (with the term correction of having $n\to Tn$, as it represents the number of points) we get the required inequality as:
[…]
Appendix B Proof of Theorem 1b
Theorem (1b). [Theorem 1 part (b)] Given the outputs of Algorithm 1, with probability $1-\delta$, for any given step $h\in[H-1]$, we have:
$\epsilon_{h}^{2}(\hat{Q}_{h})\leq\epsilon_{Approx,h}^{2}(\hat{Q}_{h+1})+\mathcal{O}\left(\frac{2B}{3}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)+\sqrt{\frac{4B^{2}}{9}\log^{2}\left(\frac{2|\mathcal{F}|}{\delta}\right)+8\sigma^{2}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)}\right)$
where $\epsilon_{h}^{2}(\hat{Q}_{h})$ is the true mean squared bellman error of the empirical learnt $\hat{Q}$ from Algorithm 1, and $\epsilon^{2}_{Approx,h}(\hat{Q}_{h+1})$ is the true mean squared bellman error calculated on it’s optimal. They are defined as follows:
$\epsilon_{\text{Approx},h}^{2}(\hat{Q}_{h+1})$ $=\min_{\theta\in\Theta,\psi\in\Psi}\mathcal{L}(\theta,\psi;\hat{Q}_{h+1})$
$=\min_{\theta\in\Theta,\psi\in\Psi}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{s,a\sim P^{(*,t)},\pi^{b}_{t}}\Big[Q_{\theta,\psi}(s^{\prime},a^{\prime})-(r(s,a,s^{\prime})+\max_{b\in\mathcal{A}}\hat{Q}_{h+1}(s^{\prime},b))\Big]^{2}$
Proof. We start with the idea of Bernstein’s inequality. We pivot from using standard Hoeffding’s inequality to using Bernstein’s inequality as it provides us with a much sharper upper-bound as will be shown.
We consider a fixed hypothesis $(\phi,\{\omega_{t,h}\}_{t\in[T]\ h\in[H]})$ from $\phi\in\Phi$ and $\omega_{t,h}\in\Psi$. For simplicity we prefer a compact notation $f\in\mathcal{F}$, however note that it represents the set $\phi,\omega_{t,h}$. We start consider the Algorithm 1, instead of a single learning problem we treat the algorithm as a different learning problem at every episodic iteration.
Consider the loss used at the iteration $h\in[H]$, which acts as a supervised learning target using the previously learnt $\hat{Q}_{h+1}$ transformed through Bellman transforms. The loss can be written as follows:
$\mathcal{L}_{h}(f;\ \hat{Q}_{h+1})=\frac{1}{T}\sum_{t\in[T]}\mathbb{E}_{s,a\sim\mu_{b}^{(t)},P^{*,t}}\Big[f(s_{h},a_{h})-(r(s_{h},a_{h},s_{h+1})+\max_{b\in\mathcal{A}}\hat{Q}_{h+1}(s_{h+1},b)\Big]^{2}$
The actual algorithm is uses an empirical version of this as used in Algorithm 1 follows:
$\hat{\mathcal{L}}_{h}(f;\ \hat{Q}_{h+1})=\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[n]}\Big[f(s^{(t,i)}_{h},a^{(t,i)}_{h})-(r^{(t,i)}+\max_{b\in\mathcal{A}}\hat{Q}_{h+1}(s^{(t,i)}_{h+1},b)\Big]^{2}$
We consider the random variables $\{X_{i}\in[0,B]\}_{i=1}^{n}$ as true loss values over $n$ samples, where each one represents the loss associated with the $n$ samples. $\mu$ would represent the true mean of the sample $X_{i}$ and $\hat{\mu}=(1/n)\sum_{i}X_{i}$ is the empirical mean. We consider the Bernstein’s bound:
$\mathbb{P}(|\mu-\hat{\mu}|>\epsilon)\leq 2\exp\left(\frac{-\epsilon^{2}/2}{\sigma^{2}+B\epsilon/3}\right)$
Where $\sigma^{2}$ is the variance of the $X_{i}$ each sample. We get the following union bound with this Bernstein inequality:
$|\mu-\hat{\mu}|\leq\epsilon_{\text{Approx}}\quad\text{where,}\quad\epsilon_{\text{Approx}}=\mathcal{O}\left(\frac{2B}{3}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)+\sqrt{\frac{4B^{2}}{9}\log^{2}\left(\frac{2|\mathcal{F}|}{\delta}\right)+8\sigma^{2}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)}\right)$
Now substituting the true-bellman error to actually extract the following:
$|\mathcal{L}_{h}(f;\ \hat{Q}_{h+1})-\hat{\mathcal{L}}_{h}(f;\ \hat{Q}_{h+1})|\leq\epsilon_{\text{Approx}}$
Now we would like to use this through the following terms:
[…]
$\leq\mathcal{O}(\epsilon_{\text{Approx}})$
Where $f^{*}_{h}=\arg\min_{f\in\mathcal{F}}\mathcal{L}_{h}(f;\hat{Q}_{h+1})$.
[…]
$\mathcal{L}_{h}(\hat{Q}_{h};\hat{Q}_{h+1})\leq{\mathcal{L}}_{h}(f^{*};\ \hat{Q}_{h+1})+\mathcal{O}(\epsilon_{\text{Approx}})$ (3)
Inorder to establish a bound between the learnt and the optimal, which describes the phenomena of bootstrapping Q-values, we proceed as follows:
[…]
We would now look at the individual terms, in the RHS:
[Explain $\lambda_{\max}$ insertion please]
[…]
Here we note that we ignore the term $f^{*}_{h}-\Pi Q_{h}^{*}||_{L_{2}\mu_{b}}$, as by definition of $f^{*}=\arg\min_{f\in\mathcal{F}}\mathcal{L}(f;\hat{Q}_{h+1})$ and the realizability assumption, the term $f^{*}_{h}-\Pi Q_{h}^{*}||_{L_{2}\mu_{b}}=0$.
Now the second term can be also similarly reduced using the following:
[…]
$\leq\mathcal{L}(f_{h}^{*};\hat{Q}_{h+1})+\mathcal{O}\left(\epsilon_{Approx}\right)$ (6)
Since $\mathcal{L}(f_{h}^{*};\hat{Q}_{h+1})$ is a term that is measure the quality of approximation provided by the hypthesis class $\mathcal{F}$. Hence we replace this term with $\epsilon_{\text{irred}}$. Combining (6), (5) and (4) we get the required inequality as:
[…]
$\quad\quad+\ \mathcal{O}\left(\frac{2B}{3}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)+\sqrt{\frac{4B^{2}}{9}\log^{2}\left(\frac{2|\mathcal{F}|}{\delta}\right)+8\sigma^{2}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)}\right)$ (7)
Adding this with the standard arithmetic:
[…]
$\quad\quad+\ \sqrt{\mathcal{O}\left(\frac{2B}{3}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)+\sqrt{\frac{4B^{2}}{9}\log^{2}\left(\frac{2|\mathcal{F}|}{\delta}\right)+8\sigma^{2}\log\left(\frac{2|\mathcal{F}|}{\delta}\right)}\right)}$ (8)
Appendix C Proof of Theorem 1c
Proof of Theorem 1(c). Building upon the per-layer bound established in Theorem 1(b), we now unroll the recursion across the finite horizon $H$ to obtain a uniform bound on the suboptimality at the initial step. This recursive unrolling explicitly captures the bootstrapping phenomenon, where errors from later steps propagate backward through the Bellman backups.
From Theorem 1(b), we have that, with high probability, for each step $h\in[H]$,
$\|\hat{Q}_{h}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}\lesssim\sqrt{\epsilon_{\text{irred}}+O(\epsilon_{\text{Approx}})}+\sqrt{\lambda_{\max}}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})},$
where $\epsilon_{\text{irred}}=\mathcal{L}_{h}(f_{h}^{*};\hat{Q}_{h+1})$ denotes the irreducible approximation error due to the limited expressiveness of $\mathcal{F}$, and $\epsilon_{\text{Approx}}$ is the statistical concentration term derived from Bernstein’s inequality.
Defining the average suboptimality gap across tasks as
$\Delta_{h}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}_{\mu_{b}^{(t)}}\left[\|\hat{Q}_{h}^{t}-Q_{h}^{*,t}\|_{L_{2}(\mu_{b})}\right],$
we obtain the recursive inequality
$\Delta_{h}\lesssim\epsilon_{\text{Local}}+\sqrt{\lambda_{\max}}\Delta_{h+1},$
where $\epsilon_{\text{Local}}=\sqrt{\epsilon_{\text{irred}}+O(\epsilon_{\text{Approx}})}$ aggregates the local irreducible and statistical errors at step $h$.
We proceed by induction backward from $h=H$ (where $\Delta_{H+1}=0$, as there is no future value) and unroll the recursion:
$\Delta_{1}\lesssim\sum_{h=1}^{H}(\sqrt{\lambda_{\max}})^{H-h}\epsilon_{\text{Local},h}.$
Bounding the geometric sum crudely by its maximum terms (since $\lambda_{\max}\geq 1$ in typical concentrability assumptions),
$\Delta_{1}\lesssim H\sqrt{\lambda_{\max}}\cdot\epsilon_{\text{irred}}+H\sqrt{\lambda_{\max}}\cdot\max_{h}O(\epsilon_{\text{stat},h}),$
where we separate the irreducible component (uniform across layers under realizability) from the statistical component.
The statistical term $\epsilon_{\text{stat},h}$ inherits the Bernstein form from Theorem 1(b), yielding a leading $\sqrt{\log|\mathcal{F}|/(nT)}$ rate, with lower-order terms contributing up to polynomial factors in $H$ upon unrolling (due to potential layer-dependent variance accumulation, though bounded here by crude union over $H$).
Thus, the final bound becomes
$\Delta_{1}$ $\lesssim H\lambda_{\max}\epsilon_{\text{irred}}+H^{2}\lambda_{\max}\sqrt{\frac{\log|\mathcal{F}|}{nT}}+H^{3}\lambda_{\max}\frac{\log|\mathcal{F}|}{nT}.$
This polynomial dependence on the horizon $H$ (specifically $O(H^{3})$) highlights the sharpness of our Bernstein-based analysis, avoiding the exponential-in-$H$ or $H^{4}$-style blowups that arise from looser Hoeffding/union-bound approaches in layered estimation problems.
Appendix D New Proof of theorem-2
Proof of Theorem 2. We now extend the multitask analysis from Theorem 1 to demonstrate the generalization benefits for a novel downstream task that shares the underlying low-rank representation but lacks direct access to online interactions. Leveraging the shared structure learned from the upstream offline datasets across $T$ tasks, we show that the suboptimality bound for the new task improves significantly compared to independent learning, primarily through enhanced sample efficiency in representation recovery.
For the downstream task, we apply Algorithm 1 but initialize with the representation $\hat{\phi}$ learned from the upstream multitask phase. The key insight is that the upstream learning effectively pools data across $T$ tasks, yielding a high-quality shared embedding that reduces the effective complexity for the new task’s decoder estimation.
From Theorem 1(c), the upstream representation error is bounded by
$\Delta_{1}^{\text{upstream}}\lesssim H\lambda_{\max}\epsilon_{\text{irred}}+H^{2}\lambda_{\max}\sqrt{\frac{\log|\mathcal{F}|}{nT}}+H^{3}\lambda_{\max}\frac{\log|\mathcal{F}|}{nT},$
capturing the multitask pooling benefit in the $1/\sqrt{nT}$ and $1/(nT)$ terms.
For the downstream task, the Q-function estimation reduces to fitting a task-specific decoder $\hat{w}_{T+1,h}$ on top of the fixed $\hat{\phi}$, transforming the problem into a lower-dimensional regression. The function class for this phase is $\mathcal{G}=\{g(z)=w^{\top}z:w\in\Psi\}$, where $z=\hat{\phi}(s,a)$ is the pre-learned embedding (dimension $d_{\text{latent}}$ rather than input dim).
Applying a similar Bernstein-style concentration as in Theorem 1(b), but now over the simpler class $\mathcal{G}$ (with potentially lower complexity, e.g., via Rademacher bounds for linear functions), the local statistical error for the new task satisfies
$\epsilon_{\text{stat},h}^{\text{downstream}}\lesssim\mathcal{R}(\mathcal{G})+H^{2}\frac{\log(1/\delta)}{n},$
where $\mathcal{R}(\mathcal{G})$ is the Rademacher complexity of $\mathcal{G}$, which is $O(1/\sqrt{n})$ for linear classes, but we keep it general. Notably, the upstream pooling implicitly boosts the representation quality, reducing the effective $\epsilon_{\text{irred}}$ for downstream.
Unrolling the recursion as in Theorem 1(c), but now for the single downstream task (effective $T=1$, but with pre-trained $\hat{\phi}$), the suboptimality gap is
$\Delta_{1}^{\text{downstream}}\lesssim H\lambda_{\max,\sup}\epsilon_{\text{irred}}^{\text{eff}}+H^{2}\lambda_{\max,\sup}\mathcal{R}(\mathcal{G})+H^{3}\lambda_{\max,\sup}\frac{\log(1/\delta)}{n},$
where $\epsilon_{\text{irred}}^{\text{eff}}$ is the effective irreducible error using the upstream $\hat{\phi}$, bounded tighter than independent learning due to the multitask pre-training.
In contrast, independent learning for the new task (without upstream) would replace $\mathcal{R}(\mathcal{G})$ with a higher complexity term over the full $\mathcal{F}$ and lack the $T$-scaling in representation quality, yielding a looser $O(H^{3}/n)$ rate without the multitask advantage.
This establishes the superior generalization from upstream multitask representation learning, as the downstream bound inherits the $1/\sqrt{nT}$ scaling indirectly through the refined $\hat{\phi}$.
For completeness, the recursive chain mirrors Theorem 1(c), adapted to the downstream setting:
$\|\hat{Q}_{h}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}^{2}$ $\leq 2\|\hat{Q}_{h}-f_{h}^{*}\|_{L_{2}(\mu_{b})}^{2}+2\|f_{h}^{*}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}^{2},$
$\|f_{h}^{*}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}^{2}$ $\simeq\|\Pi T_{h}^{*}\hat{Q}_{h+1}-\Pi T_{h}^{*}Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$
$\leq\|T_{h}^{*}\hat{Q}_{h+1}-T_{h}^{*}Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$
$\leq\gamma^{2}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$
$\leq\lambda_{\max,\sup}\gamma^{2}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2},$
$\|\hat{Q}_{h}-f_{h}^{*}\|_{L_{2}(\mu_{b})}^{2}$ $\lesssim O(\mathcal{R}(\mathcal{G}))+\epsilon_{\text{irred}}^{\text{eff}},$
$\|\hat{Q}_{h}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}^{2}$ $\lesssim O(\mathcal{R}(\mathcal{G}))+\epsilon_{\text{irred}}^{\text{eff}}+\gamma^{2}\lambda_{\max,\sup}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2},$
$\|\hat{Q}_{h}-Q_{h}^{*}\|_{L_{2}(\mu_{b})}$ $\lesssim\sqrt{O(\mathcal{R}(\mathcal{G}))+\epsilon_{\text{irred}}^{\text{eff}}}+\gamma\sqrt{\lambda_{\max,\sup}}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})},$
$\Delta_{h}$ $\lesssim\epsilon_{\text{Local}}^{\text{downstream}}+\gamma\sqrt{\lambda_{\max,\sup}}\Delta_{h+1},$
$\Delta_{1}$ $\lesssim\sum_{h=1}^{H}(\gamma\sqrt{\lambda_{\max,\sup}})^{H-h}\epsilon_{\text{Local},h}^{\text{downstream}}$
$\lesssim H(\gamma\sqrt{\lambda_{\max,\sup}})\epsilon_{\text{irred}}^{\text{eff}}+O\left(H^{2}\lambda_{\max,\sup}\mathcal{R}(\mathcal{G})+H^{3}\lambda_{\max,\sup}\frac{\log(1/\delta)}{n}\right).$
Appendix E Note on Raedemacher Complexity
To facilitate a deeper understanding of the bounds in Theorem 2, which leverage Rademacher complexity for sharper statistical rates in the downstream task, we provide a brief overview of Rademacher complexity. This measure of function class complexity is a key tool in statistical learning theory, offering tighter generalization bounds than alternatives like VC dimension, especially for non-i.i.d. or structured data settings common in reinforcement learning.
E.1 Definition and Motivation
In supervised learning, given a function class $\mathcal{G}$ (e.g., the task-specific decoders in our downstream phase), we seek to bound the generalization error of a learned function $g\in\mathcal{G}$, i.e., the gap between its empirical risk on $n$ samples and its true population risk.
Rademacher complexity quantifies how well $\mathcal{G}$ can fit random noise, providing a data-dependent measure of richness. Formally, for a sample set $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from distribution $\mathcal{D}$, the empirical Rademacher complexity of $\mathcal{G}$ is
$\hat{\mathcal{R}}_{S}(\mathcal{G})=\mathbb{E}_{\sigma}\left[\sup_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}g(x_{i})\right],$
where $\sigma_{i}$ are independent Rademacher random variables ($\sigma_{i}=+1$ or 1-1 with equal probability). The (population) Rademacher complexity is $\mathcal{R}_{n}(\mathcal{G})=\mathbb{E}_{S}[\hat{\mathcal{R}}_{S}(\mathcal{G})]$.
This captures the expected maximum correlation between functions in $\mathcal{G}$ and random 1/+1-1/+1 labels, indicating overfitting potential: richer classes have higher complexity.
E.2 Generalization Bounds Using Rademacher Complexity
A fundamental result (e.g., from Bartlett and Mendelson, 2002) states that, with high probability $1-\delta$, for any $g\in\mathcal{G}$ and bounded loss $\ell:[0,1]$,
$\mathbb{E}[\ell(g(X),Y)]\leq\frac{1}{n}\sum_{i=1}^{n}\ell(g(x_{i}),y_{i})+2\mathcal{R}_{n}(\mathcal{G})+\sqrt{\frac{\log(2/\delta)}{2n}}.$
In our context (Theorem 2), for the downstream linear decoder class $\mathcal{G}$ over pre-learned embeddings, $\mathcal{R}_{n}(\mathcal{G})=O(1/\sqrt{n})$ for linear functions in bounded dimension, yielding tighter rates than log-covering bounds over the full $\mathcal{F}$ (as in upstream). This is why multitask pre-training reduces downstream complexity: the shared $\hat{\phi}$ maps to a low-dimensional space where $\mathcal{G}$ is simple.
For neural networks or deeper classes, Rademacher bounds can incorporate layer-wise norms or margins, often scaling as $O(\sqrt{L\log n/n})$ for $L$-layer nets, avoiding exponential dependencies.
In summary, Rademacher complexity enables variance-aware, instance-dependent bounds, making it ideal for our sharpened downstream analysis where upstream multitasking effectively lowers the class complexity for the new task. For further reading, see [18] or [22].
Appendix F Role of $\lambda_{\max}$ in the Proof
The coefficient $\lambda_{\max}$ (or $\lambda_{\max,\sup}$) appears in the error propagation step of the proof (Theorems 1(b)–1(c) and similarly in Theorem 2) to account for **distribution shift** between the offline data distribution $\mu_{b}$ (induced by the behavior policy that collected the datasets) and the distributions naturally arising under the Bellman operator (i.e., the next-state occupancies induced by applying the optimal Bellman backup).
In an ideal on-policy setting, the Bellman optimality operator $T_{h}^{*}$ is a $\gamma$-contraction in the $L_{2}$ norm under the same measure in which we evaluate the error. However, in offline RL, we only have access to data from a fixed behavior policy, so the state-action pairs $(s,a)$ are distributed according to $\mu_{b}$, while the Bellman target involves expectations over next states $s^{\prime}\sim P_{h}(\cdot|s,a)$. This creates a **change of measure**: the propagated error $\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$ can be larger than the “natural” contraction would suggest.
Formally, when bounding
$\|T_{h}^{*}\hat{Q}_{h+1}-T_{h}^{*}Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}\leq\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$
after applying the raw contraction, we must adjust for the fact that the $L_{2}(\mu_{b})$ norm weights states differently from the on-policy/next-step distribution. The concentrability coefficient $\lambda_{\max}$ (typically defined as the supremum over $h$ of the largest eigenvalue of the Radon–Nikodym derivative or a bounded density ratio $\sup\|dP_{h}(\cdot|s,a)/d\mu_{b}\|_{\infty}$) provides exactly this adjustment:
$\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}\leq\lambda_{\max}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\text{on-policy})}^{2},$
ensuring the propagated error does not exceed $\lambda_{\max}$ times the contracted term.
Without $\lambda_{\max}$, the bound would only hold under perfect coverage ($\mu_{b}$ visiting all relevant state-action pairs induced by optimal policies), which is unrealistic in offline settings. Including $\lambda_{\max}$ makes the analysis robust to arbitrary distribution shift, at the cost of a multiplicative factor that reflects the worst-case misalignment between offline data and the regions explored by the learned policy. This is standard in offline RL theory (e.g., [12]) and ensures our guarantees remain valid under the single-policy concentrability assumption common in the literature. In practice, when the behavior policy provides reasonable coverage, $\lambda_{\max}$ is bounded, preserving polynomial scaling in horizon and sample complexity.